OpenStudy (anonymous):

can you show me step by step and what rule do I use in this calculus problem and why? using d/dx (1 + x)/ |/x\|)=

zepdrix (zepdrix):

I can't quite read what you've written. Is this the problem? \[\Large\bf\sf \frac{d}{dx}\frac{1+x}{|x|}\]

OpenStudy (anonymous):

The equation is surrounded by an open and close parenthesis and for the bottom the x is square root.

myininaya (myininaya):

\[\frac{d}{dx} \frac{1+x}{\sqrt{x}} ?\]

OpenStudy (anonymous):

Yes, with ( ) surrounding the entire problem.

myininaya (myininaya):

You can separate the fraction and then write each term in the form x^n then use the power rule for both terms to differentiate

OpenStudy (anonymous):

Can you please show me?

OpenStudy (anonymous):

\[ \frac{a+b}{c}=\frac ac +\frac bc \]

OpenStudy (anonymous):

@jim_thompson5910 @JA1 @Mertsj @Ashleyisakitty please show me this problem step by step please

OpenStudy (anonymous):

@JA1 can you show me